OpenStudy (anonymous):

Consider the function f(x)=2/5x-4

OpenStudy (anonymous):

a.) Find the inverse of f(x) and name it g(x). Show and explain your work. b.) Tell why or how you know that f(x) and g(x) are inverses of each other. c.) Draw the graphs of f(x) and g(x) on the same coordinate plane. Explain what about your graph shows that the functions are inverses of each other.

OpenStudy (solomonzelman):

Okay, so your function is: \(\large\color{black}{ f(x)=\frac{\LARGE 2}{\LARGE 5}x-4 }\) is that correct?

OpenStudy (anonymous):

Yes that's correct

OpenStudy (solomonzelman):

okay, so for your first step: Replace f(x), with y, and you get. \(\large\color{black}{ y=\frac{\LARGE 2}{\LARGE 5}x-4 }\)

OpenStudy (solomonzelman):

Then, switch the x and y. Like put x in place of y, and put y in place of y.

OpenStudy (solomonzelman):

please write what you get... (take your time if you need to)

OpenStudy (solomonzelman):

I mean put in y in place of x, (not in place of y)

OpenStudy (anonymous):

\[x = \frac{ 2 }{ 5 } y - 4\]

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

then solve for y.

OpenStudy (solomonzelman):

add 4 to both sides (write what you get after this)

OpenStudy (anonymous):

x = \[4 \frac{ 2 }{ 5 } y\]

OpenStudy (solomonzelman):

\(\large\color{black}{ x+4=y\frac{\LARGE 2}{\LARGE 5} }\) like this.

OpenStudy (anonymous):

Ohhh I didn't add 4 to the x

OpenStudy (solomonzelman):

then multiply every thing times 5/2.

OpenStudy (solomonzelman):

what happens to \(\large\color{black}{y\frac{\LARGE 2}{\LARGE 5} }\) when you multiply it times \(\large\color{black}{\frac{\LARGE 5}{\LARGE 2} }\) ? what happens to \(\large\color{black}{x }\) when you multiply it times \(\large\color{black}{\frac{\LARGE 5}{\LARGE 2} }\) ? what happens to \(\large\color{black}{4 }\) when you multiply it times \(\large\color{black}{\frac{\LARGE 5}{\LARGE 2} }\) ? answer these question for me, what does each thing change to?

OpenStudy (anonymous):

I know for #2 it would be 5/2x right?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

for number 1, wouldn't 2/5 cancel with 5/2 ?

OpenStudy (anonymous):

You'd be left with just y wouldn't you?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

I got disconnected sorry for late reply. and how about number 3?

OpenStudy (anonymous):

So 4 would be 4/1, so you'd end up after reducing, with 10/1, right?

OpenStudy (solomonzelman):

yes, it will be 10.

OpenStudy (solomonzelman):

So when you multiply \(\large\color{black}{ x+4=y\frac{\LARGE 2}{\LARGE 5} }\) times 5/2, you get: \(\large\color{black}{ \frac{\LARGE 5}{\LARGE 2} x+10=y }\)

OpenStudy (solomonzelman):

so you inverse function is \(\large\color{black}{f^{-1}(x)= \frac{\LARGE 5}{\LARGE 2} x+10 }\)

OpenStudy (anonymous):

So that inverse function would be my g(x), for part A of the problem

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

Now to know that your inverse function is actually the inverse function, you can: 1) Graph each function, and see if they are symmetrical about a line y=x. 2) Show that: \(\large\color{black}{f^\circ g(x)=g^\circ f(x) }\)

OpenStudy (solomonzelman):

I would assume that you need to do the second thing most likely.

OpenStudy (anonymous):

Part C is actually draw the graphs of f(x) and g(x) on the same coordinate plane. Explain what about your graph shows that the functions are inverses of each other.

OpenStudy (solomonzelman):

you have: \(\large\color{black}{f(x)= \frac{\LARGE 2}{\LARGE 5} x-4 }\) \(\large\color{black}{g(x)= \frac{\LARGE 5}{\LARGE 2} x+10 }\) are you good with part B, sure?

OpenStudy (anonymous):

Yes I am. Thank you very much :). You're a life saver

OpenStudy (solomonzelman):

Lol, you are lowering my status I am a Champion jk

OpenStudy (anonymous):

Haha, that you are!

OpenStudy (solomonzelman):

okay, so lets go with part C. this is the other way (1st one that I posted) to show that f(x) and g(x) are inverse. basically graphing them. I'll graph them for you. https://www.desmos.com/calculator/orkagazgha the \(\normalsize\color{red}{\rm red}\) is f(x) the \(\normalsize\color{blue}{\rm blue}\) is g(x) (the inverse) the the \(\normalsize\color{green}{\rm green}\) is y=x.

OpenStudy (solomonzelman):

you see that the graphs of the \(\normalsize\color{red}{\rm f(x)}\) and \(\normalsize\color{blue}{\rm g(x)}\) a reflections (mirrors) of each other across the \(\normalsize\color{green}{\rm y=x}\).

OpenStudy (anonymous):

So, g(x) being a reflection of f(x), would show that, on a graph they're inverses of each other?

OpenStudy (solomonzelman):

not exactly, g(x) being a reflection of f(x) \(\normalsize\color{black}{\rm \underline{across~~the~~line~~y=x}}\), would show that, on a graph they're inverses of each other?

OpenStudy (solomonzelman):

I edited your statement to be true.

OpenStudy (anonymous):

Ohhh, okay that makes more sense now

OpenStudy (solomonzelman):

If you have any questions, ask...

OpenStudy (anonymous):

Okay :). I think I'm all good for now. Thank you so much

OpenStudy (solomonzelman):

Anytime:)